OpenStudy (anonymous):

subistution method 7x+2y=0 x=40-6y

OpenStudy (katrinakaif):

7(40 - 6y) + 2y = 0 Distribute the 7 280 - 42y + 2y = 0 280 - 40y = 0 Bring 280 to the other side by substracting it -40y = -280 Divdie by -40 and you should get your answer for y. Then plug it into any one of the otrginal equations to get your x value.